A U.S.-based importer, Zarb Inc., makes a purchase of crystal glassware from a firm in

Switzerland for 39,960 Swiss francs, or $24,000, at the spot rate of 1.665 francs per dollar. The terms

of the purchase are net 90 days, and the U.S. firm wants to cover this trade payable with a forward

market hedge to eliminate its exchange rate risk. Suppose the firm completes a forward hedge at the

90-day forward rate of 1.682 francs. If the spot rate in 90 days is actually 1.6380 francs, how much will

the U.S. firm have saved or lost in U.S. dollars by hedging its exchange rate exposure? Enter your answer

rounded to two decimal places. Do not enter $ or comma in the answer box. For example, if your answer is

$12,300.456 then enter as 12300.46 in the answer box.
Answer #1

Solution:-

Spot rate: SF1.665/$1

Purchase price: SF39,960

Thus, SF39,960 was the forex exposure that Zarb inc. hedged

Purchase price of forex in USD on the date of hedging transaction= 39,960/1.682= $23,757.43

Thus the firm locked in its dollar cost of forex that it will require in 90-days at $23,757.43

Now, if the firm hadn't entered into the forward agreement, the dollar cost it would have had to incurr after 90 days is as follows:

Dollar cost of forex that would have been incurred after 90 days if hedging wasn't done= 39,960/1.638= $24,395.60

Thus, dollar profit on hedging= $24,395.60-$23,757.43= 638.17